If a retailer sells a coat for $139.99. What was his markup if he bought the coat for $81.79? (15 pts)

• Markup in actual dollars?

• Express markup as a % of cost?

• Express markup as a % of selling price?

*Remember that markup percentage is based on selling price unless otherwise noted.
Answer #1

Markup in Actual dollars = 139.99-81.79 = $58.2

Markup on cost = 58.2/81.79x100 = 71.15%

Markup on selling price = 58.2/139.99x100 = 41.57%


answered by: ANURANJAN SARSAM